# FormulaQA: A Question Answering Dataset for Formula-Based Numerical Reasoning

### v2

@@ 5.2 Results of Fine-tuned Small Models
5.3 Results of LLMs with Formula Retriever
Model id test (zero-shot) id test (few-shot) GPT-3.5-turbo 26.13 32.07 GPT-3.5-turbo w/ FR 22.09 33.02 GLM4 65.32 62.47 GLM4 w/ FR 70.31 65.80 Qwen-max 58.67 58.91 Qwen-max w/ FR 56.06 58.91
Table 6: Results of LLMs with Formula Retriever on FormulaQA. “w/ FR” in the table means that the model is evaluated with top-5 retrieved formulas. Prompts can be found in Appendix A.2.4.
The results of LLMs experiments utilizing the formula retriever are shown in Table 6. We found that the impact on performance varied among different LLMs when incorporating retrieved formulas into prompts. We observed a positive enhancement on GLM4, reaching the highest performance among all models, with scores of 70.31 and 65.80 on the id test with zero-shot and id test with few-shot, respectively. This indicated that incorporating a formula retriever is indeed effective, but there is still ample room to explore how to integrate the capabilities of formula retrievers with other LLMs.
5.4 Case Study and Error Analysis
@@
We sampled 50 error cases from the id test (few-shot setting) of GPT-3.5-turbo and manually categorized the types and proportions of errors. We divided the error types into two main categories: formula errors and calculation errors. Formula errors encompass inappropriate formulas and omitted formulas, while calculation errors primarily involve inaccuracies in numerical calculation and unit errors. We found that 38% of errors were caused by incorrect formulas, while the remaining 62% were attributable to calculation errors. We provide one example for each of the two types of errors listed in Figure 2. It could be observed that FormulaQA poses challenges to existing models in terms of formula application and numerical calculation (including unit calculation and arithmetic calculation).
6 Conclusion
In this paper, we introduced FormulaQA, a question answering dataset for formula-based numerical reasoning, which consists of 5,420 questions collected from Chinese junior high school physics exams. Unlike previous datasets where only a fraction of the questions relied on commonsense formulas or lacked formula annotations, we annotated the reasoning steps with formulas for each question with both manual and LLM-assisted efforts. Furthermore, we constructed a formula database by merging formulas with similar meanings, serving as an external knowledge base for subsequent retrieval-based/augmented approaches.  We evaluated FormulaQA across various sizes of LLMs, fine-tuned small models, and retrieval-augmented LLMs, demonstrating its challenging nature as an unresolved task. Additionally, FormulaQA introduces numerical reasoning questions based on domain-specific formulas, a dimension scarcely addressed in prior datasets. Our findings indicate substantial room for improvement of existing models on formula-based numerical reasoning, thus motivating future research efforts.
Limitations
In order to inspire further research, we have also translated the dataset into English unitizing LLMs. However, we have not yet accurately assessed the quality of the translated dataset. At present, we have not released the English version of the dataset, but we will do so later after ensuring the quality of the English dataset. Additionally, our dataset is limited to the domain of physics. Although junior high school physics is not overly complex and can be understood by most people, it is still possible to explore formula-based question answering data in other domains.
Ethical Considerations
We collected examination questions from publicly available data sources, which are openly accessible and drawn from examinations across various regions in China. Our dataset focuses on the elementary physics domain and its development may contribute to natural language reasoning, automated AI systems, and physics education. We do not foresee any potential risks for now. Annotators recruited for this task were undergraduate students majoring in science and engineering from a university in China and they are all skilled in elementary physics. All annotators were compensated with course hours and academic credits.
Appendix A Appendix
A.1 Dataset Construction
A.1.1 Prompts in Formula Normalization.
The process of formula normalization is delineated into three distinct stages: the generation of natural language explanations, the extraction of the associated parameters from the explanations, and the subsequent error correction phase. The initial two stages are illustrated in Figures 3 and 4. The third stage is further splited into three specific error categories, each addressed by a dedicated prompt: input errors, where the parameters mentioned in the explanation are absent from the question; calculation errors, which occur when the calculator reports an error during the computation process; and output errors, where the final computed answer is incorrect. We provide an example here focusing on prompts for correcting calculation errors, while prompts for the other two error types can be found in our code submission. The prompts designed to correcting calculation errors are depicted in Figure 5. The entire normalization procedure employs a 6-shot prompting, an instance of which is provided herein for illustrative purposes.
A.1.2 Examples of deleted questions
A.1.3 Semantics Based Merging for Formula Database Construction
Semantics-based merging primarily employs the LLM to comprehend formulas, ascertain if two formulas are semantically equivalent, and subsequently determine whether they can be merged into a single formula. The prompt for this procedure is illustrated in Figure 7. This approach ensures that the nuanced meanings embedded within formulas are accurately captured and evaluated for potential merging, thereby enhancing the quality of formula database.
A.2 Experiments
A.2.1 Zero-shot and Few-shot Prompts
@@ A.2.2 Formula Retriever
A.2.3 Few-shot Experiments on the LLMs of Chat Versions
Model Size id test ood test Avg. zero-shot CoTs with LLMs of chat versions InternLM-chat 20B 5.70 4.60 5.17 Qwen-chat 14B 32.07 37.60 34.73 Qwen-chat 7B 7.36 8.70 8.01 InternLM-chat 7B 7.84 7.67 7.76 Llama2-chat 7B 0.00 0.26 0.13 few-shot CoTs with LLMs of base versions InternLM-base 20B 18.29 11.25 14.90 Qwen-base 14B 44.89 36.83 41.01 Qwen-base 7B 21.14 18.16 19.71 InternLM-base 7B 9.50 8.18 8.86 Llama2-base 7B 1.43 0.26 0.87 few-shot CoTs with LLMs of chat versions InternLM-chat 20B 11.58 10.10 10.87 Qwen-chat 14B 41.38 37.93 39.72 Qwen-chat 7B 19.21 15.02 17.19 InternLM-chat 7B 10.10 7.88 9.03 Llama2-chat 7B 1.97 1.00 1.50
Table 7: Results of different versions of the LLMs with zero-shot and few-shot on FormulaQA.
In this experiment, we could compare the performance of the same version of the model under zero-shot and few-shot settings. Results are shown in Table 7. For the chat version of the LLMs, we could observe that few-shot can effectively improve model performance, with performance improvements ranging from 1.27 to 9.18 on average across id test and ood test. Comparing the performance of the base version and chat version of the same model under few-shot settings, except for minimal changes on InternLM-chat-7B and Llama2-chat-7B, the performance of the other models showed a decrease.
A.2.4 Prompts for LLMs with Formula Retriever
@@
Figure 7: Prompt for semantics-based merging.

### v3

@@ A.2 Case Study and Error Analysis
We sampled 50 error cases from the id test (few-shot setting) of GPT-3.5-turbo and manually categorized the types and proportions of errors. We divided the error types into two main categories: formula errors and calculation errors. Formula errors encompass inappropriate formulas and omitted formulas, while calculation errors primarily involve inaccuracies in numerical calculation and unit errors. We found that 38% of errors were caused by incorrect formulas, while the remaining 62% were attributable to calculation errors. We provide one example for each of the two types of errors listed in Figure 2. It could be observed that FormulaReasoning poses challenges to existing models in terms of formula application and numerical calculation (including unit calculation and arithmetic calculation).
A.3 Experiments
A.3.1 Data Augmentation (DA) for FormulaReasoning
There have been several studies utilizing large language models (LLMs) for data augmentation (Ding et al., 2024). The data generated in these related works (Zheng et al., 2023; Whitehouse et al., 2023) primarily focus on daily conversations or sentiment analysis and do not require rigorous numerical calculations. Some research on data augmentation involving numerical calculations (Shum et al., 2023) employs LLMs to generate solutions to questions to aid in training, rather than creating complete questions. In contrast to these approaches, our work generates complete questions that involve numerical calculations (particularly formula calculations), along with automatic improvement and selection to ensure data quality.
In order to enhance the capabilities of models, we use LLM to generate more data for fine-tuning. We divide the process of data generation into the following several steps.
First, we randomly generated 17,000 prompts. Each prompt was obtained by stacking five question-answer pairs sampled form training set. At the end of the prompt, LLM was required to generate the sixth question-answer pair. Second, we normalized the generated formulas. Except for the absence of manual review, the remaining steps were consistent with those in Section 3.2. At last, we unitized the calculator to check whether the calculation process in the data generated by the LLM is correct, and discarded the generated data with incorrect calculation processes. After the above steps, we finally retained more than 2500 questions.
We found that mixing the newly generated data into the original training set did not always bring positive improvement, perhaps because the newly generated data has not undergone manual review. We found that randomly selecting a small portion of the newly generated data can enable the model to have performance improvement. We set several different mixing ratios selected from $\{5\%,10\%,15\%,20\%,2\%,30\%,35\%,40\%\}$. We fine-tuned the ChatGLM-6B-base using the augmented data set. After training for a fixed number of steps (150k and 200k), we selected the checkpoints with the smallest loss among models of different mixing ratios.
A.4 Implementation Details
We accessed to GPT-4-turbo, GPT-3.5-turbo [^5] https://platform.openai.com/docs, GLM4666https://open.bigmodel.cn/, and Qwen-max [^7] https://help.aliyun.com/zh/dashscope/developer-reference/quick-start through API calls with the default hyper-parameters. For other LLMs, we conducted experiments on NVIDIA V100-32G GPUs for 7B models, and on NVIDIA A100-80G GPUs for 14B/20B models. These LLMs generated using nucleus sampling with top_p=0.8. Models that require fine-tuning were experimented on NVIDIA V100 GPUs with Huggingface Transformers and Pytorch 2.0. For mT5-base and mT5-large, we set a learning rate of 5e-5 and a batch size of 32, testing the model after training for 50 epochs. For Qwen-1.8B, we used a learning rate of 1e-5 and a batch size of 32, and tested the model after training for 10 epochs. For ChatGLM3-6B, we fine-tuned with LoRA Hu et al. (2021) with r=8, alpha=32 and learning rate of 5e-5, batch size of 1. The max input length and output length are both set to 512. We utilized nucleus sampling with top_p=0.8 for generation. In the case of CoT-SFT, which directly outputted formulas along with corresponding parameter values and units, if the generation output contained formatting errors, we allowed the small model to retry up to 5 times until a correctly formatted output was generated. Training mT5-base, mT5-large, Qwen-1.8B, ChatGLM-6B models requires 6, 12, 12 and 24 hours respectively.
A.4.1 Zero-shot and Few-shot Prompts
@@ A.4.2 Formula Retriever
Models were evaluated with top-5 retrieved formulas. Prompts can be found in Appendix A.4.4. We utilized zero-shot CoTs.
A.4.3 Few-shot Experiments on the LLMs of Chat Versions
Model Size id test ood test Avg. zero-shot CoTs with LLMs of chat/instruct versions InternLM-chat 20B 5.70 4.60 5.17 Qwen-chat 14B 32.07 37.60 34.73 Llama3-instruct 8B 22.66 17.98 20.41 Llama2-chat 7B 0.00 0.26 0.13 Qwen-chat 7B 7.36 8.70 8.01 InternLM-chat 7B 7.84 7.67 7.76 few-shot CoTs with LLMs of base versions InternLM-base 20B 18.29 11.25 14.90 Qwen-base 14B 44.89 36.83 41.01 Llama3-base 8B 12.81 8.87 10.91 Llama2-base 7B 1.43 0.26 0.87 Qwen-base 7B 21.14 18.16 19.71 InternLM-base 7B 9.50 8.18 8.86 few-shot CoTs with LLMs of chat/instruct versions InternLM-chat 20B 11.58 10.10 10.87 Qwen-chat 14B 41.38 37.93 39.72 Llama3-instruct 8B 6.90 6.16 6.54 Llama2-chat 7B 1.97 1.00 1.50 Qwen-chat 7B 19.21 15.02 17.19 InternLM-chat 7B 10.10 7.88 9.03
Table 7: Results of different versions of the LLMs with zero-shot and few-shot on FormulaReasoning.
In this experiment, we compared the performance of the same version of the model under zero-shot and few-shot settings. Results are shown in Table 7. For the chat version of the LLMs, we could observe that few-shot can effectively improve model performance, with performance improvements ranging from 1.27 to 9.18 on average across id test and ood test. Comparing the performance of the base version and chat version of the same model under few-shot settings, except for minimal changes on InternLM-chat-7B and Llama2-chat-7B, the performance of the other models showed a decrease from base to chat versions.
A.4.4 Prompts for LLMs with Formula Retriever
@@
Figure 7: Prompt for semantic-based merging.